OpenStudy (anonymous):

Express the complex number in trigonometric form. 3 - 3i

OpenStudy (anonymous):

\[put3-3\iota=r \cos \alpha+\iota r \sin \alpha\] equate real and imaginary parts and calculate r and \[\alpha \]

OpenStudy (loser66):

take a look at this, for yours, just put everything in formula as surjithayer said. on my paper, I use \(\theta\) and R instead of r and \(\alpha\) and r as his. Everything else is the same http://openstudy.com/study#/updates/5206789de4b01dab33708fe7